OpenStudy (anonymous):

1.A diprotic acid weighing 6.5 grams is neutralized by 137.5 mL of 0.75M NaOH. What is the molar mass? 2. 20,000 gallons of nitric acid (72 % acid) density is 1.42g/cc. How many liters of 6M Sodium Bicarbonate is needed to neutralize? @matt101

OpenStudy (cuanchi):

If is is neutralized that means that the number of H from the acid is equal to the number of OH from the base. Calculate how many moles of NaOH you have in 137.5mL of 0.75M n= M x V (L) number of moles equal to molarity times the volume in liters Because the acid is diprotic divide that value by 2 to calculate how many moles of the acid you have in the 6.5 g Then calculate the molecular mass of the acid n= m/MM number of moles is equal to the mass in g divided the molecular mass, then the MM, will be the mass (6.5g) divided the number of moles of acid calculated in the previous step.

OpenStudy (cuanchi):

NaHCO3 + HNO3→ NaNO3 + H2O + CO2 1) calculate the moles of Nitric acid in the 20,000 gallons (72% v/v) 1.42g/cc convert gallon to liters (1gal = 3.78541 L) calculate the 72% of this value and convert the L to cc then to g with the value of density (1.42g/cc) convert the g to moles (n=m/MM) MM nitric acid =63.01 g/mol 2) convert the moles of Nitric acid to volume of sodium bicarbonate n= M x V(L) => V(L) = n/M

OpenStudy (matt101):

Looks like you got a good answer. Let us know if you have any questions @heathred!

OpenStudy (anonymous):

Okay so for the first one, I would calculate moles which would be (0.1375L)(0.75M)=0.103125 then I would divide that by 2 making it 0.0515625. Then how do I get the MM? @Cuanchi

OpenStudy (matt101):

You have moles and you have mass. How do you get MM from that?

OpenStudy (anonymous):

Well since moles=mass/molar mass, wouldn't I multiply moles and mass?

OpenStudy (matt101):

Not quite - rearrange the equation to solve for molar mass and what do you get?

OpenStudy (anonymous):

mass divided by moles?

OpenStudy (matt101):

Yup!

OpenStudy (anonymous):

I am still confused with the second question? When does the 6M come into play?

OpenStudy (matt101):

You'll first need to find the moles of sodium bicarbonate needed to neutralize the nitric acid solution. The question gives you the concentration of the bicarbonate solution (6M), so using that and the moles you calculated will allow you to find the volume!